OpenStudy (anonymous):

Can someone help me find the LCM of 49y^2+245y and 7y^2+56y+105

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

49y and 7y

OpenStudy (mathmate):

The best way to find the LCM is to factorize each one: 49y^2+245y=49y(y+5) 7y^2+56y+105=7(y^2+8y+15)=7(y+5)(y+3) So the LCM is 49y(y+5)(y+3), i.e. without repeating the common factors.

OpenStudy (anonymous):

mathmate do you think you could help with one more that has me stumped

OpenStudy (mathmate):

Just did!